OpenStudy (anonymous):

Verify Identity 1-2secx-3(sec^2)x/-(tan^2)x=1-3secx/1-secx

OpenStudy (sidsiddhartha):

tan²(x) = sec²(x) - 1=(secx-1)(secx+1), again (1-2secx-3sec^2x)=(1 - 3sec(x))(1 + sec(x)) then it will be =-(1 - 3sec(x))(1 + sec(x))/((sec(x) - 1)(sec(x) + 1)) = -(1 - 3sec(x))/(sec(x) - 1) = RHS